# Axi-majoron for almost everything

### v1

@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
where we used $\Gamma^{\phi}\approx\Gamma_{a}^{\phi}$ in the second line. Since axi-majorons produced by misalignment are the only contribution to dark matter and kinetic misalignment does not work in our scenario (see Section 5), it is required to have $\phi_{0}\sim\phi_{0}^{\rm ref}$. Also, the validity of Eq. (4.1) is limited to the case of $\phi_{*}\lesssim\phi_{\times}$ which puts a limit $\xi_{\phi}\lesssim 100$ (see Eq. (3.70)). Moreover, the constraint of $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ (see Eq. (3.82)) requires ${\rm B}_{1}\gtrsim\left(m_{\nu_{1}}/m_{\nu}\right)^{1/2}$. Therefore, if $c_{\rm AD}$ is not much smaller than unity, $Y_{\Phi,e}$ should be smaller than unity by a couple of orders of magnitude in order to match observation.
If $\phi_{*}>\phi_{\times}$, $\Phi$ field decays before falling into the region of the true vacuum position. As shown in Eq. (3.63), this case would be possible if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ for $\phi_{0}\sim\phi_{0}^{\rm ref}$. Thermalization of $N_{1}$ does not occur as long as $\xi_{\phi}\sim\mathcal{O}(10^{2})$ while keeping $\phi_{0}\approx\phi_{0}^{\rm ref}$. The late time baryon number asymmetry in this case is given by
$Y_{B,\rm AD}^{>}$ $=$ $-\beta_{1}\times\frac{12}{37}\left(\frac{3\sqrt{3}}{2\pi}\right)^{1/2}\left(\frac{g_{*}^{\Phi}}{g_{*}(T_{\rm d})}\right)^{1/4}\lambda_{\phi}^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}Y_{\Phi,e}$ (4.13)
$\simeq$ $-7.0\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{200}{\xi_{\phi}}\right)^{1/6}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{2/3}\left(\frac{Y_{\Phi,e}}{10^{-6}}\right)$
Compared to Eq. (4.1), Eq. (4.13) does not have the suppression factor $\Delta_{\rm AD}$, and a smaller $Y_{\Phi,e}$ is required for a given set of parameters.
@@
Figure 9: Parameter for a baryon number asymmetry (red solid line), matching observations $Y_{B}^{\rm obs}\simeq 9\times 10^{-11}$ [62], from the contribution of the lightest RHN via AD-mechanism as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.1$, and $\phi_{0}=\phi_{0}^{\rm ref}$. For a smaller $Y_{\Phi,e}$, the red line is pushed up. The blue shaded region may be excluded due to too much dark radiation. Blue liens corresponds to $\Gamma_{1}/H_{\rm eq}=10^{-3}$(dotted blue), $10^{-2}$(dashed blue), and $0.1$(blue). Green lines are for $T_{\gamma,N_{1}}/m_{N_{1}}=0.05$(dashed green) and $0.1$(solid blue). Left: $Y_{B,\rm AD}^{<}$ with $Y_{\Phi,e}=10^{-1}$ in Eq. (4.1) for $\xi_{\phi}=10$ leading to $\phi_{*}/\phi_{\times}=9.1\times 10^{-3}$. Right: $Y_{B,\rm AD}^{>}$ with $Y_{\Phi,e}=10^{-6}$ in Eq. (4.13) for $\xi_{\phi}=200$ leading to $\phi_{*}/\phi_{\times}=3.6$.
@@
A remark is in order before closing this subsection. The initial asymmetry stored in $\Phi$ is only partially transferred to the visible sector with the help of RHNs. The remaining part of the asymmetry is kept as the background motion of axi-majoron. Since $U(1)_{\rm PQ}$ symmetry is broken only spontaneously in the matter sector, the amount of the net PQ-number asymmetry in a co-moving volume after inflation should be conserved. This means that the loss of asymmetry in the visible sector should appear as the motion of axi-majoron, but has minor impact on the overall homogeneous motion. Meanwhile, if the background axi-majoron motion is large enough, it can become the main source of baryogenesis via spontaneous leptogenesis as discussed in Ref. [61] for example. We discuss this possibility in the next subsection.
@@ 4.2 Spontaneous leptogenesis via axi-majoron
If RHNs are in thermal bath while they are coupled to the derivative of the axi-majoron field, the idea of spontaneous baryogenesis [40] can be applied. In Ref. [61], it was shown that in the presence of a non-zero $\dot{\theta}$, the derivative couplings of the majoron(the axi-majoron in our scenario) to fermion currents can act as a source term for generating non-zero lepton number asymmetry via inverse decay processes ($\ell+H\to N_{i}^{\pm}$) in equilibrium through the Yukawa interaction $y_{\nu}$. In our scenario, due to its nature of sharing the properties of axion and majoron, axi-majoron contributes to QCD sphaleron [63] as well as electroweak sphaleron processes, and Boltzmann equations are modified similarly to the case discussed in Ref. [39]. However, once $L$-number violation is frozen when the inverse decay to $N_{i}^{\pm}$ is closed, $B-L$ is also frozen. Also, $B+L$ is eventually driven to cancel axi-majoron contribution in the electroweak sphaleron contribution in Boltzmann equations, and $\dot{\theta}/T$ at the freeze-out epoch of the sphalerons is negligible in our scenario as will be clear in the subsequent discussion. Hence, the eventual baryon and lepton number asymmetries are determined by the non-zero $B-L$ number which is determined in a way similar to Ref. [61]. The effect of QCD/electroweak sphalerons on the results is minor, and we will ignore such a difference since it is easily compensated by other parameters.
If $N_{1}$ decays out of equilibrium and is never thermalized that is the case discussed in the previous section, only heavy RHNs are relevant for spontaneous leptogenesis. The temperature of the SM thermal bath established by their decays is given in Eq. (3.86), and one finds
@@
with $x_{\rm fo}\approx 10$. In our scenario $H_{\rm fo}$ is larger than the one related to $T_{\gamma,\rm fo}$ in the standard cosmology, leading to a smaller $x_{\rm fo}$. However, such a dependence is mainly logarithmic on ${\rm B}_{N}$, and minor as long ${\rm B}_{N}=\mathcal{O}(0.1)$ which we assume in this work. In this reason, we ignore such a dependence.
@@
Hence, $\dot{\theta}T_{\gamma}^{2}\propto a^{-5}$ for ${\tilde{\phi}}_{*}\ll\phi_{0}$. Since the standard thermal bath is established by the decay of the lightest RHN $N_{1}$ which should dominate the universe for a while, there would be a dilution of pre-existing baryon and lepton number asymmetries. If $N_{1}$ decays well before the epoch of the EWPT, the late time baryon number asymmetry through the spontaneous baryogenesis mechanism is given by [64]
@@
where $Y_{B-L}^{\rm d}$ is the yield of $B-L$ charge at the epoch of $N_{1}$’s decay. One finds
$Y_{B-L}^{\rm d}$ $=$ $Y_{B-L}^{\rm fo}\left(\frac{s_{\rm fo}}{s_{\rm d}}\right)\left(\frac{a_{\rm fo}}{a_{\rm d}}\right)^{3}$ (4.21)
$=$ $4{\rm B}_{N}x_{\rm fo}Y_{B-L}^{\rm fo}\left(\frac{T_{\rm d}}{m_{N}}\right)\left(\frac{H_{\rm fo}}{H_{\rm eq}}\right)^{1/2}$
where $T_{\rm d}$ is the decay temperature of $N_{1}$, and $H_{\rm eq}\equiv H_{1,\rm eq}$ given by Eq. (3.81). Note that Eq. (4.21) is valid for $H_{\rm fo}/H_{\rm eq}\geq 1$. The $B-L$ asymmetry at freeze-out is given by
$Y_{B-L}^{\rm fo}=-\frac{45c_{B-L}}{12\pi^{2}g_{*S}(T_{\gamma,\rm fo})}\left(\frac{\dot{\theta}}{T}\right)_{\rm fo}\ ,$ (4.22)
where $c_{B-L}=\mathcal{O}(1)$ [61]. In the case of $\phi_{*}<\phi_{\times}$, the velocity of axi-majoron at the epoch of $H_{\rm fo}$ is given by
$\dot{\theta}_{\rm fo}\approx\dot{\theta}_{e}\left(\frac{\phi_{\times}}{\phi_{e}}\right)\left(\frac{H_{*}}{H_{\times}}\right)^{2}\left(\frac{H_{\rm fo}}{H_{*}}\right)^{3/2}\ ,$ (4.23)
with
$H_{\times}\equiv\frac{\lambda_{\phi}^{1/2}\phi_{\times}^{2}}{2\sqrt{3}M_{\rm P}}\ .$ (4.24)
Also, the decay temperature of $N_{1}$ is found as
$\frac{T_{\rm d}}{m_{N}}=\left(\frac{5}{2\pi^{4}g_{*}(T_{\rm d})}\right)^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{b_{1}}{b_{N}}\right)^{1/2},$ (4.25)
and
$\left(\frac{H_{\rm fo}}{H_{\rm eq}}\right)^{1/2}$ $=$ $c_{\rm inv}^{1/2}\left(\frac{b_{N}m_{\nu_{N}}}{b_{1}m_{\nu_{1}}}\right)^{1/2}\left(\frac{\Gamma_{N_{1}}}{H_{\rm eq}}\right)^{1/2}$ (4.26)
$=$ $\left[\frac{3c_{\rm inv}}{4\sqrt{2}}\frac{b_{N}}{b_{1}{\rm B}_{1}^{2}\lambda_{\phi}^{1/2}}\left(\frac{m_{\nu_{N}}\phi_{0}}{v_{h}^{2}}\right)\right]^{1/2}\ .$
@@
$Y_{B,\rm sb}^{<}$$\approx$ $-\frac{2c_{B-L}c_{\rm inv}^{2}x_{\rm fo}^{2}}{\pi^{9/2}g_{*}(T_{\rm d})^{1/4}g_{*}(T_{\gamma,\rm fo})}\frac{b_{N}^{2}\lambda_{\phi}^{5/4}}{b_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{h}^{2}}\right)^{2}Y_{\Phi,e}$ (4.27)
$\simeq$ $-3.1\times 10^{-10}\left(\frac{0.05}{b_{1}}\right)\left(\frac{b_{N}}{0.15}\right)^{2}\left(\frac{\xi_{\phi}}{15}\right)^{5/2}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}Y_{\Phi,e}\ .$
@@
where we used $c_{B-L}=2,\ g_{*}(T_{\rm d})=g_{*}(T_{\gamma,\rm fo})=106.75$, and $m_{\nu_{N}}=m_{\nu}$ in the second line. Note that, interestingly, $Y_{B}$ does not depend on the symmetry-breaking scale.
Figure 10: Parameter space of $H_{\rm fo}/H_{1,\rm eq}=1$(red line), $H_{\rm d}/H_{1,\rm eq}\leq 0.1$(blue solid line), $0.01$(blue dashed line) with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$, and $\phi_{*}/\phi_{\times}=1$(green line) for $b_{1}=0.05,\ b_{2}=b_{3}=0.15$, and $m_{\nu_{N}}=m_{\nu}=10^{3}m_{\nu_{1}}$. The position of green line is nearly independent with branching ratios of $N_{i}$s. In the region left of the green line, since $\Gamma_{N_{i}}/H_{1,\rm eq}\propto b_{i}\phi_{0}/{\rm B}_{1}^{2}b_{1}\xi_{\phi}$, reducing $b_{2,3}$ shifts the blue line upward while red lines are also move upward due to the $b_{N}$-dependence of ${\rm B}_{1}$ even for a fixed $b_{1}$. A smaller $m_{\nu_{1}}$ also pushes up the blue lines.
In Fig. 10, we depicted parameter spaces for different cases. The condition $H_{\rm fo}/H_{\rm eq}\geq 1$ is satisfied in the region left of the red line. The blue region is nearly excluded due to too much dark radiation which is discussed in Section 5. However, the region can be pushed up by taking $m_{\nu_{1}}$ small enough. We also show in Fig. 11 the parameter space for $Y_{B,\rm sb}^{<}$ matching observations with $\Gamma_{1}/H_{\rm eq}\ll 1$ satisfied to recover the standard thermal bath. In the figure, as $\xi_{\phi}$ increases, the red line moves down or right while blue lines and shaded region move to opposite direction, as clear from Eqs. (3.82) and (4.27).
Figure 11: The late time baryon number asymmetry, $Y_{B,\rm sb}^{<}/Y_{\Phi,e}=Y_{B}^{\rm obs}$(red line), in Eq. (4.27) as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.15$, $\xi_{\phi}=15$, and $\phi_{0}=\phi_{0}^{\rm ref}$. The scheme of the blue region and the blue lines is the same as the one in Fig. 9. In the region left of the dashed green line, $H_{\rm fo}>H_{\rm eq}$ is satisfied.
As shown in Fig. 10, and inferred from Eq. (4.26), if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$, the case of $H_{\rm fo}<H_{\rm eq}$ can also arise. In such case the heavy RHN states can also dominate the universe, and recover the standard thermal bath by their decay (see Eq. (3.82)). It allows the possibility of dark matter made of mainly the lightest RHNs when $\phi_{0}$ is not large enough to saturate the relic density by axi-majoron dark matter. Keeping this possibility in mind, we consider the case of $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ with all RHNs having nearly same mass parameters and Yukawa couplings $y_{\nu}$. In regard of baryogenesis, this case is nearly the same as the case of the lightest stable and two nearly degenerate heavy RHN states, as long as the branching fraction to the lightest one is sufficiently small in order for it not to over-close the universe. Note that, as can be seen from Eq. (3.63), a large $\xi_{\phi}$ also leads to $\phi_{*}>\phi_{\times}$, i.e., the universe is dominated by a radiation-like energy of $\phi$-field at the epoch of $\phi$-particles’ decay. The expansion rate at the epoch of the inflaton’s decay in this case is then given by
$H_{*}=\frac{\lambda_{\phi}^{1/2}}{2\sqrt{3}}\frac{\phi_{*}^{2}}{M_{\rm P}}\ .$ (4.28)
Since RHNs having nearly the same decay rates are responsible for the final reheating,
$\frac{H_{\rm fo}}{H_{\rm d}}\approx\frac{3}{2}c_{\rm inv}\ ,$ (4.29)
where $H_{\rm d}\equiv(2/3)\Gamma_{N}$, and the late time baryon number asymmetry is given by
$Y_{B,\rm sb}^{>}=\frac{12}{37}Y_{B-L}^{\rm fo}$ (4.30)
for $N_{i}$s decaying well before the epoch of the EWPT. The velocity of axi-majoron at the epoch of $H_{*}$ is given by

### v2

@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
where we used $c_{\rm AD}=0.1,\ \Gamma^{\phi}\approx\Gamma_{a}^{\phi}$, and $g_{*,\rm d}=106.75$ in the second line. Since axi-majorons produced by misalignment are the only contribution to dark matter and kinetic misalignment does not work in our scenario (see Section 5), it is required to have $\phi_{0}\sim\phi_{0}^{\rm ref}$. Also, the validity of Eq. (4.15) is limited to the case of $\phi_{*}<\phi_{\times}$ which puts a limit $\xi_{\phi}\lesssim 100$ (see Eq. (3.71)). Moreover, the constraint of $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ requires ${\rm B}_{1}\gtrsim\left(m_{\nu_{1}}/m_{\nu}\right)^{1/2}$ (see Eq. (3.83)). Therefore, if $c_{\rm AD}$ is not much smaller than unity, $Y_{\Phi,e}$ should be smaller than unity by a couple of orders of magnitude in order to match observation.
If $\phi_{*}>\phi_{\times}$, the speed of $N_{1}$ at its decay is then given by
$\beta_{1,\rm d}^{>}=\frac{m_{\phi,*}}{2m_{N_{1}}}\left(\frac{H_{1,\rm eq}}{H_{*}}\right)^{1/2}\left(\frac{H_{\rm d}}{H_{1,\rm eq}}\right)^{2/3}$ (4.16)
with $H_{*}$ and $H_{1,\rm eq}$ given in Section 3.3. The yield of the helicity asymmetry of $N_{1}$ at the epoch in this case is given by
$Y_{1,\rm d}^{>}=\left(\frac{H_{*}}{H_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,*}}Y_{\Phi,e}\ .$ (4.17)
Note that, compared to Eq. (4.14), Eq. (4.17) does not have the suppression factor $\Delta_{\rm AD}$. As a result, the late time baryon number asymmetry is given by
$Y_{B,\rm AD}^{>}$ $=$ $\frac{12}{37}\times\beta_{1,\rm d}^{>}Y_{1,\rm d}^{>}$ (4.18)
$\simeq$ $6.2\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi}}{200}\right)^{7/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{Y_{\Phi,e}}{10^{-6}}\right)\ .$
Matching observation in this case requires a much smaller $Y_{\Phi,e}$ relative to the case of $\phi_{*}<\phi_{\times}$. It is notable that Eq. (4.18) does not depend on $\phi_{0}$ although the relic density of dark matter requires $\phi_{0}\sim\phi_{0}^{\rm ref}$.
@@
Figure 10: Parameter for a baryon number asymmetry (red solid line), matching observations $Y_{B}^{\rm obs}\simeq 9\times 10^{-11}$ [71], from the contribution of the lightest RHN via AD-mechanism as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.1$, and $\phi_{0}=\phi_{0}^{\rm ref}$. Left: $Y_{B,\rm AD}^{<}$ with $Y_{\Phi,e}=10^{-1}$ in Eq. (4.15) for $\xi_{\phi}=10$ leading to $\phi_{*}/\phi_{\times}=9.1\times 10^{-3}$. Right: $Y_{B,\rm AD}^{>}$ with $Y_{\Phi,e}=10^{-6}$ in Eq. (4.18) for $\xi_{\phi}=200$ leading to $\phi_{*}/\phi_{\times}=3.6$. For a smaller $Y_{\Phi,e}$, the red line is pushed up. The blue shaded region may be excluded due to too much dark radiation. Blue lines corresponds to $\Delta N_{\rm eff}=0.1$(dotted), $0.2$(dashed), and $0.5$(solid). Green lines are for $T_{\rm d}/m_{N_{1}}=0.02$(dotted), $0.05$(dashed), and $0.1$(solid).
@@
A remark is in order before closing this subsection. The initial asymmetry stored in $\Phi$ is only partially transferred to the visible sector with the help of RHNs. The remaining part of the asymmetry is kept as the background motion of axi-majoron. Since $U(1)_{\rm PQ}$ symmetry is broken only spontaneously in the matter sector, the amount of the net PQ-number asymmetry in a co-moving volume after inflation should be conserved. This means that the loss of asymmetry in the visible sector should appear as the motion of axi-majoron, but has minor impact on the overall homogeneous motion. Meanwhile, if the background axi-majoron motion is large enough, it might become a large source of baryon number asymmetry via spontaneous leptogenesis as discussed in Refs. [41, 72, 69] for example. We discuss this possibility in the next subsection.
@@ 4.2 Spontaneous leptogenesis via axi-majoron
If RHNs are in thermal bath while they are coupled to the derivative of the axi-majoron field, the idea of spontaneous baryogenesis [42] can be applied. In Ref. [69], it was shown that in the presence of a non-zero $\dot{\theta}$, the derivative couplings of the majoron(the axi-majoron in our scenario) to fermion currents can act as a source term for generating non-zero lepton number asymmetry via inverse decay processes ($\ell+H\to N_{i}^{\pm}$) in equilibrium through the Yukawa interaction $y_{\nu}$. In our scenario, due to its nature of sharing the properties of axion and majoron, axi-majoron contributes to QCD sphaleron [73] as well as electroweak sphaleron processes, and Boltzmann equations are to be modified correspondingly. However, once $L$-number violation is frozen when the inverse decay to $N_{i}^{\pm}$ is closed, $B-L$ is also frozen. Also, $B+L$ is eventually driven to cancel axi-majoron contribution in the electroweak sphaleron contribution in Boltzmann equations, and $\dot{\theta}/T$ at the freeze-out epoch of the sphalerons is negligible in our scenario as will be clear in the subsequent discussion. Hence, the eventual baryon and lepton number asymmetries are determined by the non-zero $B-L$ number which is determined in a way similar to Ref. [69]. The effect of QCD/electroweak sphalerons on the results is minor, and we will ignore such a difference since it is easily compensated by other parameters.
As discussed in the previous section, $N_{1}$ decays out of equilibrium and is never thermalized. Hence, only heavy RHNs are relevant for spontaneous leptogenesis in our scenario. The temperature of the SM thermal bath established by their decays is given in Eq. (3.88), and one finds
@@
with $x_{\rm fo}\approx 10$. In our scenario, ${\tilde{H}}_{\rm fo}$ is larger than the one related to $T_{\gamma,\rm fo}$ in the standard cosmology, leading to a smaller $x_{\rm fo}$. However, such a dependence is mainly logarithmic on ${\rm B}_{N}$, and minor as long ${\rm B}_{N}=\mathcal{O}(0.1)$ which we assume in this work.
@@
Hence, $\dot{\theta}T_{\gamma}^{2}\propto a^{-5}$ for ${\tilde{\phi}}_{*}\ll\phi_{0}$. The standard thermal bath is expected to be established by the decay of the lightest RHN $N_{1}$ well before the EWPT. Since $N_{1}$s should dominate the universe for a while as non-relativistic particles, there would be a dilution of pre-existing baryon and lepton number asymmetries. The late time baryon number asymmetry through the spontaneous baryogenesis mechanism is given by
@@
where $Y_{B-L}^{\rm d}$ is the yield of $B-L$ charge at the epoch of $N_{1}$’s decay. It depends on whether the freeze-out of inverse processes producing heavy RHNs takes place before or after $N_{1}$ starts dominating the universe (i.e., whether ${\tilde{H}}_{\rm fo}\geq{\tilde{H}}_{1,\rm eq}$ or not).
For $\phi_{*}<\phi_{\times}$, one finds
$\frac{{\tilde{H}}_{\rm fo}}{{\tilde{H}}_{1,\rm eq}}$ $=$ $\frac{3c_{\rm inv}}{2}\frac{b_{N}}{b_{1}{\rm B}_{1}^{2}\lambda_{\phi}^{1/2}}\left(\frac{m_{\nu_{N}}\phi_{0}}{v_{h}^{2}}\right)$ (4.26)
$\simeq$ $1.4\times\left(\frac{b_{N}}{b_{1}}\right)\left(\frac{0.1}{{\rm B}_{1}}\right)^{2}\left(\frac{10}{\xi_{\phi}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$
Figure 11: Parameter space of ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}=1$(orange dashed line), $\Delta N_{\rm eff}=0.5$(blue solid line), $0.2$(blue dashed line), $0.1$(blue dotted line), and $\phi_{*}/\phi_{\times}=1$(purple dashed line) for $b_{1}=0.05,\ b_{N}=0.1$, and $m_{\nu_{N}}=m_{\nu}=10^{4}m_{\nu_{1}}$. The black dotted line is the upper bound of $\xi_{\phi}$ given in Eq. (3.70). The position of purple line does not depend on $b_{i}$s and $m_{\nu_{i}}$s. A larger/smaller $b_{1}$ with other parameters fixed shifts the orange dashed line to the left/right (see Eq. (4.26)). The blue shaded regions and blue lines are move in the same way, but a smaller $m_{\nu_{1}}$ pushes them to the left.
Fig. 11 shows the parameter spaces for ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}$. In the region of ${\tilde{H}}_{\rm fo}\geq{\tilde{H}}_{1,\rm eq}$, one finds
$Y_{B-L}^{\rm d}$ $=$ $\left(\frac{s_{\rm fo}}{s_{\rm d}}\right)\left(\frac{a_{\rm fo}}{a_{\rm d}}\right)^{3}Y_{B-L}^{\rm fo}$ (4.27)
$=$ $4{\rm B}_{N}x_{\rm fo}\left(\frac{T_{\rm d}}{m_{N}}\right)\left(\frac{{\tilde{H}}_{\rm fo}}{{\tilde{H}}_{1,\rm eq}}\right)^{1/2}Y_{B-L}^{\rm fo}\ .$
The $B-L$ asymmetry at freeze-out is given by
$Y_{B-L}^{\rm fo}=\frac{45c_{B-L}}{12\pi^{2}g_{*S}(T_{\gamma,\rm fo})}\left(\frac{\dot{\theta}}{T}\right)_{\rm fo}\ ,$ (4.28)
where $c_{B-L}=\mathcal{O}(1)$ [69]. The velocity of axi-majoron at the epoch of ${\tilde{H}}_{\rm fo}$ is given by
$\dot{\theta}_{\rm fo}\approx\dot{\theta}_{e}\left(\frac{\phi_{\times}}{\phi_{e}}\right)\left(\frac{{\tilde{H}}_{*}}{H_{\times}}\right)^{2}\left(\frac{{\tilde{H}}_{\rm fo}}{{\tilde{H}}_{*}}\right)^{3/2}\ .$ (4.29)
@@
$Y_{B,\rm sb}^{<}$$\approx$ $\frac{3.3c_{B-L}c_{\rm inv}^{2}x_{\rm fo}^{2}}{\pi^{9/2}g_{*,\rm d}^{1/4}g_{*,\rm fo}}\frac{b_{N}^{2}\lambda_{\phi}^{5/4}}{b_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{h}^{2}}\right)^{2}Y_{\Phi,e}$ (4.30)
$\simeq$ $1.5\times 10^{-10}\left(\frac{0.05}{b_{1}}\right)\left(\frac{b_{N}}{0.1}\right)^{2}\left(\frac{\xi_{\phi}}{50}\right)^{5/2}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{Y_{\Phi,e}}{0.1}\right)\ ,$
@@
where we used $c_{B-L}=2,\ g_{*,\rm d}=g_{*,\rm fo}\equiv g_{*}(T_{\rm fo})=106.75$, and $m_{\nu_{N}}=m_{\nu}$ in the second line. Note that, $Y_{B}^{<}$ does not depend on the symmetry-breaking scale.
Comparing Eq. (4.30) to Eq. (4.15), one finds
$\frac{Y_{B,\rm sb}^{<}}{Y_{B,\rm AD}^{<}}\simeq 55\times{\rm B}_{1}^{1/3}b_{3}{\rm B}_{3}\left(\frac{m_{\nu}}{10^{8}m_{\nu_{1}}}\right)^{2/3}\ .$ (4.31)
It implies that in order for the contribution from spontaneous leptogenesis to become relevant, $m_{\nu_{1}}$ should be smaller than $m_{\nu}$ by many orders of magnitude. However, $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ as shown in Eq. (4.5) and Fig. 8, and hence $Y_{B,\rm sb}^{<}$ is always subdominant, relative to $Y_{B,\rm AD}^{<}$ for a given set of parameters.
Depending on $b_{1}$, the freeze-out of the inverse processes could take place when $H={\tilde{H}}_{\rm fo}^{\prime}<{\tilde{H}}_{1,\rm eq}$ while $\phi_{*}<\phi_{\times}$ is maintained. From the scaling behavior of $Y_{B-L}$ after the freeze-out epoch, one can see that, relative to the case of ${\tilde{H}}_{\rm fo}>{\tilde{H}}_{1,\rm eq}$, $Y_{B-L}^{\rm d}$ in this case is suppressed by a factor
$\left(\frac{{\tilde{H}}_{1,\rm eq}}{{\tilde{H}}_{\rm fo}}\right)\left(\frac{{\tilde{H}}_{\rm fo}^{\prime}}{{\tilde{H}}_{1,\rm eq}}\right)^{1/2}<1\ .$ (4.32)
That is, the contribution to the baryon number asymmetry in the visible sector is always sub-dominant than the one in Eq. (4.30). Hence, the contribution from AD-mechanism is expected to be always dominant in this case, too, and we do not consider the details of this case further.
As the last possibility arising in spontaneous leptogenesis in our scenario, as shown in Fig. 11, and inferred from Eq. (4.26), if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ with $b_{1}$ not much smaller than $b_{N}$, the case of $H_{\rm fo}<H_{1,\rm eq}$ can arise with $\phi_{*}>\phi_{\times}$, too. The $B-L$ asymmetry at the epoch of $N_{1}$’s decay in this case is given by
$Y_{B-L,\rm d}^{>}$ $=$ $\left(\frac{s_{\rm fo}}{s_{\rm d}}\right)\left(\frac{a_{\rm fo}}{a_{\rm d}}\right)^{3}Y_{B-L}^{\rm fo}$ (4.33)
$=$ $4{\rm B}_{N}x_{\rm fo}\left(\frac{T_{\rm d}}{m_{N}}\right)\left(\frac{H_{\rm fo}}{H_{1,\rm eq}}\right)^{2/3}Y_{B-L}^{\rm fo}\ ,$
where $H_{1,\rm eq}$ is to be taken from Eq. (3.90). Since the freezing-out of the inverse processes in this case takes place during matter-domination era, $H_{\rm fo}$ in Eq. (4.33) is larger than the one in radiation-domination era. If $H_{1,\rm eq}>H_{\rm fo}\gg H_{\rm d}$, ignoring the contribution of the partial decay of $N_{1}$ to the SM thermal bath, one can express $H_{\rm fo}$ as
$H_{\rm fo}=H_{1,\rm eq}\left(\frac{x_{1,\rm eq}}{x_{\rm fo}}\right)^{3/2}\ ,$ (4.34)
where $x_{1,\rm eq}\equiv m_{N}/T_{1,\rm eq}$ with $T_{1,\rm eq}$ being the temperature when $N_{1}$ starts dominating the universe. In this case, $x_{1,\rm eq}$ is found as
$x_{1,\rm eq}=\left(\frac{3\pi^{2}}{80}g_{*,\rm eq}\right)^{1/4}\frac{b_{N}^{1/2}\lambda_{\phi}^{1/4}}{b_{1}^{1/2}{\rm B}_{1}{\rm B}_{N}^{1/4}}\ ,$ (4.35)
where $g_{*,\rm eq}\equiv g_{*}(T_{1,\rm eq})$. Equating the expansion rate to the interaction rate of the inverse processes, one finds that $x_{\rm fo}\sim 10$ whereas $x_{1,\rm eq}\approx 0.3-9$ for $b_{1}=0.01-1$ with $b_{N}=0.1$ as an example. Hence, considering the case as an illustrative set of parameters, we may take $x_{\rm fo}=10$ for simplicity in the subsequent discussion. The velocity of axi-majoron at the epoch of $H_{*}$ is given by